You win $100 in a basketball pool. You have


a choice between spending the money now or


putting it away for a year in a bank account that


pays 5 percent interest. What is the opportunity


cost of spending the $100 now?

1
Expert's answer
2022-03-09T10:59:01-0500

Solution

If the amount of $100 is not spent today and kept in bank account for a year then it will earn a interest of "=\frac{5}{100}\times{100} = 5"

The expected return will be;

"=(100+5)"

"=" $105


Therefore, the opportunity cost will be the difference between the expected returns of each option.

The opportunity cost of spending the $100 now

"=105-5"

"=" $5